rxcklesskaisher:

An artist is designing a kite like the one show below. Calculate the area to determine how much material she will need to create the kite. A) 38 square inches B) 180.5 square inches C) 114 square inches D) 66.5 square inches

rxcklesskaisher:

1 attachment
Florisalreadytaken:

the formula of finding the area of the kite would be \

rxcklesskaisher:

ok so divide

Florisalreadytaken:

wait, why did the formula delete😭

Florisalreadytaken:

\

rxcklesskaisher:

so

Florisalreadytaken:

something is wrong :/

rxcklesskaisher:

what

Florisalreadytaken:

1 attachment
Florisalreadytaken:

the formula is not showing up

Florisalreadytaken:

a bug or something

rxcklesskaisher:

ohh

Florisalreadytaken:

s = (xy) times 1/2*

Florisalreadytaken:

1 attachment
rxcklesskaisher:

ok and

Florisalreadytaken:

so, we'd get s = (19 x 19) times 1/2

Florisalreadytaken:

s = 361/2

rxcklesskaisher:

i was afk

Florisalreadytaken:

its orite -- are you winning at least?

rxcklesskaisher:

winning?

Florisalreadytaken:

yeah ?

rxcklesskaisher:

confued

rxcklesskaisher:

confused*

Florisalreadytaken:

ugh!

Florisalreadytaken:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Florisalreadytaken s = 361/2 \(\color{#0cbb34}{\text{End of Quote}}\) . -- this would be your answer

rxcklesskaisher:

ohhh

Florisalreadytaken:

use the calculator right here

1 attachment
rxcklesskaisher:

ik

Florisalreadytaken:

all that said, which of the options is the right one?

rxcklesskaisher:

thats aint a answer..

rxcklesskaisher:

im so dumb..

Florisalreadytaken:

ok, so whats the correct answer?

Florisalreadytaken:

@rxcklesskaisher ?

rxcklesskaisher:

afk againn

rxcklesskaisher:

ok imma stay

Florisalreadytaken:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Florisalreadytaken all that said, which of the options is the right one? \(\color{#0cbb34}{\text{End of Quote}}\) ?

rxcklesskaisher:

the correct answer i think is a

Florisalreadytaken:

seriously...

Florisalreadytaken:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Florisalreadytaken s = 361/2 \(\color{#0cbb34}{\text{End of Quote}}\) AREA = 180.5

rxcklesskaisher:

MF B DEN

Florisalreadytaken:

finally

rxcklesskaisher:

THANK YOU